# Dinkel: Testing Graph Database Engines via State-Aware Query Generation

### v2

@@ 5. Evaluation
To demonstrate the effectiveness of Dinkel in GDBMS bug detection, we evaluate it on real-world GDBMSs and seek to answer the following questions:
@@
    Can Dinkel find real bugs in widely-used and extensively-tested GDBMSs? (Section 5.2)
@@
    How do query context and graph schema contribute to the effectiveness of Dinkel? (Section 5.4)
@@ 5.1. Experimental Setup
We evaluated Dinkel on three real-world GDBMSs, Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Apache AGE (Apache AGE. Graph database optimized for fast analysis and real-time data processing., 2024). We chose them because they are popular and open-source GDBMSs supporting Cypher. Neo4j is the most popular GDBMS, according to DB-Engines Ranking (DB-Engines Ranking., 2024), and has been extensively tested by existing approaches (Hua et al., 2023; Jiang et al., 2023b; Kamm et al., 2023; Zhuang et al., 2023). RedisGraph was widely adopted by Redis (the 6th most popular DBMS (DB-Engines Ranking., 2024)) to provide functionalities of graph data querying and storage. Its fork, FalkorDB (FalkorDB. A super fast Graph Database uses GraphBLAS under the hood for its sparse adjacency matrix graph representation., 2024), is also actively developed. Apache AGE is a widely-used extension for PostgreSQL (the 4th most popular DBMS (DB-Engines Ranking., 2024)) that enables applications to manipulate graph data on the top of a relational DBMS.
We evaluate Dinkel on these GDBMSs with the latest versions. During our testing, if the code of the tested GDBMSs is updated (e.g., a new version is released), we set up new Dinkel instances to test the updated versions. Specifically, we test Neo4j starting from version 5.6.0, RedisGraph starting from version 2.12.0, and ApacheAGE starting from version 1.3.0. We clone the code of these GDBMSs from their official GitHub repositories. To demonstrate the effectiveness of Dinkel on testing black-box GDBMSs, we also evaluate Dinkel on the Enterprise version of Neo4j from 5.6.0. Each time we implement new features on Dinkel, we stop and restart the testing. Overall, the testing process lasts three months. The evaluation was performed on Ubuntu 20.04 with a 64-core AMD EPYC 7742 processor running at 2.25GHz and 256GB of RAM.
@@ 5.2. Bug Detection
Neo4j 32 32 32
RedisGraph 17 17 10
Apache AGE 11 9 9
Total 60 58 51
Table 5 shows the status of the bugs found by Dinkel. In total, Dinkel found 60 unique bugs, including 32 bugs in Neo4j, 17 in RedisGraph, and 11 in Apache AGE. Among these bugs, 58 are confirmed, and 51 are fixed. None of the 60 bugs are duplicates.
Bug Classification. We classify the bugs found by Dinkel into two categories according to their manifestation:
  • Internal errors. The tested GDBMSs unexpectedly throw exceptions or errors when processing syntactically and semantically valid queries. The error messages can indicate the inconsistency of the internal execution status (e.g., the error triggered by the query shown in Figure 6).
  • Crashes. The queries cause GDBMSs to crash due to assertion failures or memory corruption, e.g., a null-pointer dereference.
@@
GDBMS Internal Error Crash
Neo4j 32 0
RedisGraph 3 14
Apache AGE 8 3
Total 43 17
Note that Neo4j is realized in Java and implements exception handling for unexpected errors (e.g., null dereferences in Figure 6 and out-of-bound accesses in Figure 16), and thus Neo4j does not crash but posts exception information when triggering memory corruption. Therefore, we classify all the bugs Dinkel found in Neo4j as internal errors. Table 6 shows the classification results. Among the 60 Dinkel found, 43 cause internal errors, 32 of which are Neo4j bugs. Dinkel can find 17 bugs that crash RedisGraph and Apache AGE, which are implemented in C/C++. Among these 17 bugs, 8 are caused by memory corruptions, and 9 are caused by assertion failures. These results demonstrate that Dinkel is effective in finding bugs in GDBMSs.
Affected GDBMS Components. We investigated 42 of the 51 fixed bugs, where we can analyze the fix patches to identify the GDBMS components affected by the bugs accurately. The 9 fixed bugs in the Enterprise version of Neo4j are not included because developers keep the fix patches confidential. Table 7 shows the results. Among the 42 fixed bugs, 32 affect the parsers of GDBMSs, 15 affect the planners, and 10 affect the executor. Interestingly, 9 bugs can affect two components, and 4 can affect all the components. For example, the bug shown in Figure 16 can affect both the parser and planner of Neo4j. In total, 19 bugs can affect either the planner, the executor, or both. These results demonstrate that (1) Dinkel can find bugs in various GDBMS components; and (2) Dinkel can find bugs in the deep logic of GDBMS implementation, considering 45% (19/42) of bugs are related to the planners or executors.
Table 7. GDBMS components affected by bugs
@@
Neo4j 17 11 6
RedisGraph 7 4 4
Apache AGE 8 0 1
Total 32 15 11
Diversity of Cypher Features. We investigate the queries triggering the 60 bugs and analyze the distribution of the Cypher features used in these queries. All these queries are reduced. The results are shown in Figure 9. The RETURN keyword is present in most bug-triggering queries, appearing in 64% of them. This is unsurprising, as RETURN is required to appear at the end of a query containing only reading clauses (e.g., MATCH), otherwise, the query would be syntactically incorrect. MERGE (33%) and CREATE (26%) are the two clauses inducing the creation of graph properties (e.g., providing available elements for the operations of subsequent clauses). We can deduce from this that graph data is often required to be present for bugs to trigger. The remaining keywords can induce complicated data flow (e.g., WITH (22%) and UNWIND (14%)) or control flow (e.g., CALL (14%) and CASE WHEN (14%)) in the query, indicating that bugs often appear when the database system has to handle such flows. ORDER BY also appears frequently (14%), which forces GDBMSs to keep track of the manipulated graph data and perform non-trivial comparisons to ensure correct order.
Figure 9. Feature distribution of bug-triggering queries.
Data Dependencies for Triggering Bugs. To convey the query complexity Dinkel required to trigger bugs, we analyze the complexity of the 60 bug-triggering queries. The results are shown in Figure 10. Nearly half of the bugs (45%) require at least one data dependency. Specifically, 13 bug-triggering queries contain one data dependency, 8 contain two, and 6 contain three or more. For example, in Figure 3, the bug-triggering query contains 8 data dependencies, as illustrated in Figure 11. Among the 8 dependencies, 6 are involved in query context (i.e., the variables x and y assigned to a node and a relationship), and 2 are involved in graph schema (i.e., the label A). These results demonstrate that some GDBMS bugs can be triggered only when the queries contain multiple data dependencies, and Dinkel can effectively find these bugs.
@@
Figure 12. Query sizes of bug-triggering queries.
Size of Bug-Triggering Queries. Figure 12 shows the size of the 60 bug-triggering queries. 54 bugs can be triggered by queries whose size is less than 100 bytes. The bug-triggering queries shown in Figure 3 and Figure 6 are examples of such queries. As query size increases from 0 to 100 bytes, the number of triggered bugs increases almost linearly. The bug-triggering query with the largest size is 146 bytes and is shown in Figure 16. This query triggers an internal error caused by an out-of-bounds access in Neo4j.
Bug Importance. Neo4j, the GDBMS we focus on, provides Community and Enterprise versions. The bugs in the Enterprise version are critical because this version is commonly deployed on commercial applications. Among the 32 Neo4j bugs Dinkel found, 19 bugs can be triggered in both Enterprise and Community versions, and 8 bugs can be triggered in only the Enterprise version. RedisGraph and Apache AGE did not provide severity information about the reported bugs, but we noticed that most of the bugs were fixed, which indicates that the bugs are non-trivial.
Some developers express their appreciation for our effort in finding bugs in their GDBMSs. Particularly, Neo4j provides very positive feedback written in German. We translate it into English:
@@ 5.3. Query Generation
To understand the complexity and validity of the queries generated by Dinkel, we ran Dinkel on each tested GDBMS for 24 hours and collected all generated queries. Table 8 shows the statistical results.
Table 8. Queries generated by Dinkel within 24 hours
@@
Neo4j 68K 89.88% 23.68 30.16 1202.39
RedisGraph 956K 98.14% 23.42 30.15 992.41
Apache AGE 367K 81.79% 16.94 15.60 830.35
Total 1391k 93.43% 21.72 26.31 959.85
  • Gen.: the number of generated queries;
    Valid: the number of valid queries among the generated ones;
    Clause, Dep., and Size: the average number of clauses, data dependencies, and bytes over all generated queries, respectively.
Query Complexity. For the 1391K generated queries, Dinkel constructs 30.23M clauses. The average number of clauses for each query is 21.72. Additionally, on average, each query contains 26.31 data dependencies. Because of the large number of query clauses and data dependencies, the queries generated by Dinkel are typically big. The average query size is 959.85 bytes. These results demonstrate that Dinkel can effectively generate complex queries, which are large and contain multiple clauses with complicated data dependencies. The generated queries are more complex than the bug-triggering queries discussed in Section 5.2, because all the bug-triggering queries are reduced, while the generated queries inevitably contain many redundant parts (Jiang et al., 2023c; Maciver and Donaldson, 2020; Regehr et al., 2012). We find that the queries generated for Apache AGE are less complex (i.e., fewer clauses, fewer data dependencies, and smaller size), because Apache AGE supports fewer clauses, and many clauses that can induce data dependencies are missing. As a result, the relative chance of Dinkel choosing clauses (e.g., RETURN) that terminate query generation becomes much higher, and thus the queries generated for Apache AGE exhibit lower complexity.
@@
Query Validity. Among the 1391K Cypher queries generated by Dinkel, 1300K are valid. The percentage of valid queries is 93.43%. The result demonstrates that Dinkel can keep a high validity rate even when generating complex queries. We investigated the invalid queries generated by Dinkel and found that they were mainly caused by illegal arithmetic operations. Some arithmetic operations require their operands to satisfy some constraints (e.g., the divisor in a division operation must be non-zero). However, Dinkel cannot track the value of each operator in queries because the value may depend on complex expressions, whose results are difficult to calculate (e.g., hash functions). Queries failing to satisfy the constraints of operations cause semantic errors like divisions by zero.
We noticed that the validity of queries generated for Apache AGE is lower than that of the other two GDBMSs. This is caused by the fact that Apache AGE is designed to be a PostgreSQL extension and thus needs to interact with the SQL contexts of PostgreSQL. Specifically, the columns returned by Cypher queries of Apache AGE have to match the type definitions in the wrapping SQL queries used by PostgreSQL, otherwise, the SQL queries become invalid. Figure 13 shows an example, where the Cypher query is valid but the overall SQL query is invalid because the Cypher query returns two columns while the SQL query expects only one column. Such invalid cases do not affect bug detection, because the error of SQL queries occurs only after the Cypher queries are fully executed. For example, in Figure 13, even though the SQL query is invalid, it can crash Apache AGE and PostgreSQL, because the internal Cypher query triggers the bug before it returns values to the SQL query.
1SELECT * FROM cypher('graph',\$\$
2 CREATE (x) SET x.n0 = (true OR true)
3 RETURN 0 AS y, 1 AS z
4\$\$) AS (v agtype);
Figure 13. An invalid SQL query that wraps a valid Cypher query, which crashes Apache AGE and PostgreSQL.
@@ 3 RETURN 0 AS y, 1 AS z
Throughput. On average, one Dinkel instance generates 5.36 test cases per second. Specifically, at each second, one Dinkel instance generates 0.78 test cases for Neo4j, 11 test cases for RedisGraph, and 4.25 test cases for Apache AGE. To understand the bottleneck of test throughput, we further investigated the time used by Neo4j in query generation and query execution, and found that 92.79% of CPU time is occupied by Neo4j executing the generated queries, indicating that the testing throughput is determined by the performance of the tested GDBMS. We believe the current throughput is practical considering (1) GDBMS testing typically lasts for several months (Kamm et al., 2023; Jiang et al., 2023b), and thus a sufficient number of test cases can be executed; and (2) setting up multiple Dinkel instances can significantly improve the test efficiency.
5.4. Ablation Analysis
To understand the contribution of query context and graph schema for bug detection in GDBMSs, we analyze the 60 bug-triggering queries by investigating whether the generation for these queries depends on specific graph states. Specifically, we extract the data dependencies contained by each query and check whether at least one of the data dependencies is related to query context or graph schema. Table 9 shows the analysis results.
@@ 5.4. Ablation Analysis
Among the 60 bug-triggering queries, 22 depend on only query context, 1 depends on only graph schema, and 4 depend on both (e.g., the bug-triggering query shown in Figure 11). According to these results, without query context, 26 bugs cannot be found. This demonstrates that query context is important because it enables Dinkel to generate queries referencing specific nodes, relationships, or expressions constructed in internal clauses. Without graph schema, a further 5 bugs cannot be found. These bug-triggering queries require Dinkel to properly reference the latest graph labels or properties. The following shows two queries involving query context and graph schema, respectively.
@@
Table 9. Bug-triggering queries related to graph states
@@
Neo4j 32 11 1 2
RedisGraph 17 6 0 2
Apache AGE 11 5 0 0
Total 60 22 1 4
  • QC: the queries depend on only query context;
    GS: the queries depend on only graph schema;
    Both: the queries depend on two graph states.
Example 1: Query referencing query context. The query shown in Figure 14 triggers an internal error. This query uses a UNWIND clause to iterate over the elements in the array [0]+[] (i.e., [0]). Then, the query invokes toBoolean to convert the iterated element to the boolean type and return each converted element. However, when concatenating the arrays [0] and [], Neo4j incorrectly maintains the type information of the concatenated array of [0] and []. It directly assigns the type of elements in the right-hand-side array (i.e., []), which is implicit and not determined, to elements of the concatenated array. In the end, the type of the elements in the concatenated array is unexpectedly assigned as Float (while Integer is the expected type), resulting in the internal error of type mismatching. To fix this bug, when concatenating two arrays, developers make Neo4j consider the element type in both arrays.
Example 2: Query referencing graph schema. In Figure 15, the query specifies a graph pattern ()-[:A]->({x:0}), where the relationships with type A are connecting an arbitrary node to another node with property x equal to 0. The MERGE clause tries to find the nodes and relationships satisfying the pattern. If there is no such node and relationship, the MERGE clause creates new ones satisfying the patterns. After the MERGE clause, the EXISTS clause checks whether the graph pattern (:!(B&C))-[{x:0}]->() can be matched in the current graph database and returns the corresponding boolean value. The EXISTS clause references two non-existing labels, B and C, and the property x created in the MERGE clause. As the bug is triggered in the Enterprise version of Neo4j, the root cause and the corresponding fixing patch are kept confidential by Neo4j developers.
1UNWIND [0]+[] AS i
2RETURN toBoolean(i)
Figure 14. A query that references query context and triggers a Neo4j internal error—Type mismatch: expected Boolean, Integer or String but was Float (line 1, column 37 (offset: 36))
1MERGE ()-[:A]->({x:0})
2RETURN EXISTS {(:!(B&C))-[{x:0}]->()};
Figure 15. A query that references graph schema and triggers a Neo4j internal error—Neo4jError: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.properties(…)” because ”relationshipCursor” is null).

### v3

@@ 5. Evaluation
Our evaluation aims to answer the following questions:
@@
    Can Dinkel find real bugs in widely used and extensively tested GDBMSs? (Section 5.2)
@@
    How do different techniques contribute to the effectiveness of Dinkel? (Section 5.4)
@@ 5.1. Experimental Setup
We evaluated Dinkel on Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Memgraph (Memgraph. Open-source graph database, tuned for dynamic analytics environments. Easy to adopt, scale and own., 2024), which are popular and extensively tested by existing approaches (Hua et al., 2023; Jiang et al., 2023b; Kamm et al., 2023; Zhuang et al., 2023; Mang et al., 2024b). We evaluate Dinkel on the latest GDBMS versions. During testing, if the tested GDBMSs are updated (e.g., a new version is released), we set up new Dinkel instances to test the updated versions. Specifically, we test Neo4j from version 5.6.0, RedisGraph from 2.12.0, and Memgraph from 2.7.0. RedisGraph is no longer maintained after 2.12.10, and thus we move to test its successor, FalkorDB (FalkorDB. A super fast Graph Database uses GraphBLAS under the hood for its sparse adjacency matrix graph representation., 2024), from its first release version 4.0.0. We count the bugs of RedisGraph and FalkorDB together as they share the majority of their code base. To show its effectiveness for black-box testing GDBMSs, we evaluate Dinkel on the Enterprise version of Neo4j from 5.6.0 onward. Each time we implement new features on Dinkel, we stop and restart the testing. Overall, the testing campaign intermittently lasted 9 months. The evaluation was performed on a machine running Ubuntu 20.04, with a 64-core AMD EPYC 7742 processor running at 2.25GHz and 256GB of RAM.
@@ 5.2. Bug Detection
Table 5 shows the status of the bugs found by Dinkel. In total, Dinkel found 127 unique bugs, including 51 bugs in Neo4j, 62 in RedisGraph, and 14 in Memgraph. Among these bugs, 113 are confirmed, and 84 are fixed. None of the 127 bugs are duplicates. As mature GDBMSs, Neo4j, RedisGraph, and Memgraph have been extensively tested in both industry and academia. The significant number of new bugs demonstrates the powerful bug-finding ability of Dinkel. We will publish all our 127 bug reports afterward with our artifact.
Bug Classification. We classify the bugs found by Dinkel into three categories according to their manifestation:
  • Logic bugs. The tested GDBMSs produce incorrect results for specific queries. These bugs were exposed because they lead to discrepancies between the results of the original queries and the transformed queries generated by Dinkel.
  • Internal errors. The tested GDBMSs unexpectedly throw exceptions or errors when processing syntactically and semantically valid queries. The error messages can indicate the inconsistency of the internal execution status.
  • Crashes. The queries cause GDBMSs to crash due to assertion failures or memory corruption.
@@
Neo4j 51 51 49
RedisGraph 62 48 29
Memgraph 14 14 6
Total 127 113 84
@@
GDBMS Logic Bug Internal Error Crash
Neo4j 13 38 0
RedisGraph 18 10 34
Memgraph 2 8 4
Total 33 56 38
Table 6 shows the classification results. Among the 127 bugs Dinkel found, 33 are logic bugs, 56 cause internal errors, and 38 result in GDBMS crashes. Note that Neo4j is implemented in Java and has exception handling for unexpected memory errors. Thus, Neo4j does not crash, but rather responds with exception information when triggering memory corruption. Therefore, Dinkel found no crashes in Neo4j, but 38 internal errors. Among the 38 crashes found in RedisGraph and Memgraph, 29 are caused by memory corruptions, and 9 are caused by assertion failures. These results demonstrate that Dinkel can comprehensively test GDBMSs by finding various bugs.
Figure 6 shows a bug that triggers a Neo4j internal error. This query constructs an empty array [] and invokes a subquery using a CALL. The subquery also constructs an empty array. It then invokes two UNWIND clauses, which iterate over each element in the operated array. For each iterated element, the UNWIND executes the subsequent clause under the context of this element. For example, the array [x] used by the second UNWIND references the variable x, which is 0 when the first UNWIND iterates over the item 0 in the array [0]. After the CALL, the query utilizes FOREACH, whose execution depends on the query context of the CALL. To optimize query execution, Neo4j tries to flatten the FOREACH loop. However, such optimization does not work well when the loop involves update operations (i.e., MERGE) under complicated contexts (i.e., the contexts produced by CALL). The improper optimization corrupts the internal data structures of Neo4j, Eagers, which are the production of another optimization, Eagerness analysis. In the end, an internal error is triggered when Neo4j tries to access the corrupted Eagers. To fix this bug, Neo4j developers modify both the Eagerness analysis and the flattening strategy for FOREACH clauses to ensure they work consistently.
Affected GDBMS Components. We investigated 59 of the 84 fixed bugs, where we were able to analyze the fix patches to identify the GDBMS components affected by the bugs accurately. The 19 fix patches in the Enterprise version of Neo4j are confidential and thus not included. The 6 fixed bugs in Memgraph are also not included because Memgraph developers mix the fixes into their regular development commits, where we cannot accurately analyze the affected components. Table 7 shows the results. Among the 59 fixed bugs, 29 affect the parsers of GDBMSs, 25 affect the planners, and 20 affect the executor. Interestingly, 29 bugs (5 logic bugs) affect two components, and 12 bugs (1 logic bug) affect all the components. For example, the bug shown in Figure 10 affects both the planner and executor of Neo4j. In total, 40 bugs (8 logic bugs) affect either the planner, the executor, or both. The 6 logic bugs affecting the executor also affect the planner. These results demonstrate that (1) Dinkel can find bugs in various GDBMS components; and (2) Dinkel can find bugs in the deep logic of GDBMSs, considering 67.8% (40/59) of fixed bugs are related to the planners or executors.
WITH [] AS n0 ORDER BY null
CALL {
WITH [] AS n1 ORDER BY null
UNWIND [0] AS x
UNWIND [x] AS n2
RETURN n2 AS n3
} FOREACH (n4 IN null | MERGE ({key:n3}))
Figure 6. A query triggering a Neo4j internal error—Neo4j- Error: ExecutionFailed (arraycopy: last destination index 6 out of bounds for object array[5]).
Table 7. GDBMS components affected by bugs. The number in the parentheses is for the logic bugs
@@
Neo4j 18 (0) 18 (5) 10 (3)
RedisGraph 15 (1) 15 (3) 19 (3)
Total 23 (1) 23 (8) 29 (6)
Data Dependencies for Triggering Bugs. To convey the data dependencies required to trigger bugs, we analyze dependencies in the 127 bug-triggering queries. The results are shown in Figure 7. Over half of the bugs (58%) require at least one data dependency. Specifically, 42 bug-triggering queries contain one data dependency, 16 contain two, 11 contain three, and 7 contain four or more. For example, in Figure 1, the query contains 8 data dependencies, as illustrated in Figure 8. Among the 8 dependencies, 6 are involved in query context (i.e., the variables x and y assigned to a node and a relationship), and 2 are involved in graph schema (i.e., the label A). These results demonstrate that some GDBMS bugs can be triggered only when the queries contain multiple data dependencies, and Dinkel can effectively find these bugs.
@@
Size of Bug-Triggering Queries. Figure 9 shows the size of the 127 bug-triggering queries. All these queries have been reduced using the methods mentioned in Section 4. 112 bugs can be triggered by queries whose size is less than 120 bytes. The bug-triggering query shown in Figure 1 is an example of such a query. As query size increases from 0 to 120 bytes, the number of triggered bugs increases almost linearly. The bug-triggering query with the largest size is 249 bytes, which triggers an internal error in Neo4j and was fixed by their developers.
Bug Importance. Neo4j, the GDBMS we focus on, provides Community and Enterprise versions. The bugs in the Enterprise version are critical because this version is commonly deployed on commercial applications. Among the 51 Neo4j bugs Dinkel found, 32 bugs can be triggered in both Enterprise and Community versions, and 19 bugs can be triggered in only the Enterprise version. RedisGraph and Memgraph did not provide severity information about the reported bugs, but we noticed that most bugs were fixed, indicating that the bugs are non-trivial. Some developers express their appreciation for our effort in finding bugs in their GDBMSs. Particularly, Neo4j provides very positive feedback:
@@ 5.3. Query Generation
To understand the complexity and validity of the queries constructed by Dinkel, we ran Dinkel on each tested GDBMS for 48 hours and collected all generated queries (transformed queries are not included). Table 8 shows the statistical results.
Query Complexity. On average, Dinkel generates 655k queries for each GDBMS, with 22.62M clauses constructed. The average number of clauses of each query is 34.54, and the number of data dependencies is 26.36. Because of the large number of clauses and dependencies, the queries generated by Dinkel are typically large. The average query size is 1153.3 bytes. These results demonstrate that Dinkel can effectively generate complex queries, which are large and contain multiple clauses with complicated data dependencies. The generated queries are more complex than the bug-triggering queries discussed in Section 5.2, because all the bug-triggering queries are reduced, while the generated queries inevitably contain many redundant parts (Jiang et al., 2023c; Maciver and Donaldson, 2020; Regehr et al., 2012).
Figure 9. Query sizes of bug-triggering queries.
Table 8. Queries generated by Dinkel within 48 hours. Gen.: the number of generated queries; Valid: the number of valid queries; Clause, Dep., and Size: the average of clauses, data dependencies, and bytes over all queries, respectively
@@
Neo4j 318k 89.90% 40.97 26.77 1394.69
RedisGraph 1’078k 93.90% 32.71 30.59 1071.58
Memgraph 570k 83.27% 29.95 21.71 993.63
Average 655k 89.02% 34.54 26.36 1153.3
@@
Query Validity. Among the 655k queries generated by Dinkel, 583k are valid. The percentage of valid queries is 89.02%. The result demonstrates that Dinkel can keep a high validity rate even when generating complex queries. We investigated the invalid queries generated by Dinkel and found that they were mainly caused by illegal arithmetic operations. Some arithmetic operations require the operands to satisfy some constraints (e.g., the divisor in a division must be non-zero). However, Dinkel cannot track the value of each operator in queries because the value may depend on complex expressions, whose results are difficult to calculate (e.g., hash functions). Queries failing to satisfy the constraints of operations cause semantic errors, such as divisions by zero.
@@
Throughput. On average, one Dinkel instance generates 3.79 tests (each test consists of one original query and one transformed query) per second. Specifically, at each second, one Dinkel instance completes 1.84 tests for Neo4j, 6.24 for RedisGraph, and 3.30 for Memgraph. To understand the bottleneck of test throughput, we further investigated the time used by Neo4j in query generation, transformation, and execution. We found that for Neo4j, 74.75% of CPU time was used to execute the generated queries, which are complex and can be time-consuming. This result indicates that the test throughput is dominated by the performance of the tested GDBMSs. We believe the current throughput is practical considering (1) GDBMS testing typically lasts for several months (Kamm et al., 2023; Jiang et al., 2023b), and thus a sufficient number of test cases can be executed; and (2) setting up multiple Dinkel instances can significantly improve the test efficiency.
5.4. Importance of Key Techniques
To show the importance of our techniques, we analyze how they contribute to the 127 bugs found by Dinkel. The analysis consists of two parts. The first focuses on the query context and graph schema in query generation, while the second assesses the transformation strategies (4 clause-level ones and 2 expression-level ones) used in test-oracle construction.
@@ 5.4. Importance of Key Techniques
We analyzed the 127 bug-triggering queries and checked whether their generations depend on specific graph states. Specifically, we extracted the data dependencies contained by each query and check whether at least one dependency is related to query context or graph schema. Table 9 shows the results. Among the 127 queries, 72 depend on only query context, 5 depends on only graph schema, and 9 depend on both. According to these results, without query context, 81 bugs (12 logic bugs) cannot be found. This demonstrates the importance of query contexts, which enable Dinkel to generate queries referencing specific nodes, relationships, or expressions constructed in internal clauses. Without graph schema, 14 bugs (3 logic bugs) will be missed. These bug-triggering queries require Dinkel to properly reference the latest graph labels or properties.
@@
Table 9. Bug-triggering queries related to QC: only query context; GS: only graph schema; Both: both graph states. The number in the parentheses is for the logic bugs
@@
Neo4j 51 (13) 23 (3) 3 (0) 5 (2)
RedisGraph 62 (18) 34 (7) 1 (1) 4 (0)
Memgraph 14 (2) 5 (0) 1 (0) 0 (0)
Total 127 (33) 72 (10) 5 (1) 9 (2)
Table 10. Transformation strategies used in finding logic bugs, including Dead Code (DC), Unused Context (UC), Redundant Write (RW), Inconsequential Supersession (IS), Mathematical Identity (MI), and Cypher Feature (CF)
Type Strategy Neo4j RedisGraph Memgraph Total
Clause DC 3 4 0 7
UC 2 3 0 5
RW 1 3 1 5
IS 2 6 0 8
Total 8 16 1 25
Expr MI 1 1 1 3
CF 4 1 0 5
Total 5 2 1 8
Total 13 18 2 33
We investigated the 33 logic bugs found by Dinkel, and checked whether their bug-triggering queries use specific transformation strategies. Table 10 shows the results. We find that each logic bug was triggered by using just one transformation rule in Table 3.2 and Table 3.2.1. Specifically, among the 33 logic bugs, 25 are exposed by clause-level transformations, and 8 by expression-level transformations. The number of bugs found by each transformation strategy is close to the average (i.e., 5.5). Inconsequential Supersession finds the most bugs with 8, and Mathematical Identity finds the least with 3 bugs. These results indicate that all strategies are effective in identifying logic bugs in GDBMSs. Moreover, the 25 bugs found by clause-level transformations demonstrate the effectiveness of our fine-grained transformation over EET (Jiang and Su, 2024), which supports only expression-level transformation.
The following discusses two representative bug examples.
Example 1: Figure 10 shows a Neo4j bug that causes an incorrect query result. The original query references both query contexts and graph schema. Specifically, the first CREATE clause creates two nodes with a relationship y, labeled A. One node has a property x with value 0. The second CREATE clause creates a new node with a property x whose value is equal to the property x of the end node of y (i.e., 0). The RETURN clause returns the number of nodes whose property x is 0. The expected result should be 2. However, Neo4j produces 1, because of a logic bug triggered when Neo4j processes two consecutive CREATEs that have data dependencies. Dinkel finds this bug using a clause-level transformation, Redundant Write. It inserts a CREATE clause and DELETE clause between the two CREATEs. Such a transformation preserves the query semantics but makes the two CREATEs not consecutive anymore, and thus Neo4j produces a correct result for the transformed query. The developers fixed this bug by significantly reconstructing the source code related to CREATE clauses. In the end, 17 source files were modified.
// Original: 1
CREATE ({x:0})<-[y:A]-()
CREATE ({x:endNode(y).x})
RETURN COUNT {({x:0})} AS n
// Transformed: 2
CREATE ({x:0})<-[y:A]-()
CREATE (z) DELETE z
CREATE ({x:endNode(y).x})
RETURN COUNT {({x:0})} AS n
Figure 10. Clause-level transformation reveals a bug.
Example 2: Figure 11 shows a logic bug in Neo4j found by expression-level transformation, Mathematical Identity. It is based on a rule that the result of NULL XOR a is NULL, no matter what a is. Dinkel transforms the NULL into NULL XOR EXISTS{()}, expecting that the result should remain NULL. However, Neo4j produces FALSE. The root cause of this bug is related to an operator in the Cypher planner of Neo4j, LetSelectOrSemiApply. It is invoked only when the transformed query is processed. Such an operator fails to work correctly because it does not consider NULL expressions in its arguments. To fix this bug, Neo4j developers refined this operator by adding additional logic to check and handle NULL in its arguments.
// Original: NULL
RETURN NULL AS x
// Transformed: FALSE
RETURN NULL XOR EXISTS{()} AS x
Figure 11. Expression-level transformation reveals a bug.